# Probabilistic Emulation of a Global Climate Model with Spherical DYffusion

### v1

@@ B.3 FV3GFS configuration for data generation
In the following we summarize the reference data for this study, as also discussed in Section 2.1 of [58]. The training and validation data is generated by running an ensemble of 11 10-year (after discarding a 3-month spinup period) FV3GFS simulations on a C96 cubed-sphere grid (approximately 100 km horizontal grid spacing) with 63 vertical levels. The simulations are an initial-condition ensemble. That is, they are identical except for using different initial atmospheric states. Initial-condition ensembles are a popular tool in climate modeling [26]. Discarding a 3-month spinup period ensures that each simulation is independent of each other due to the chaoticity of the atmosphere [^4] E.g. see Kay et al. [26], who note that: “After initial condition memory is lost, which occurs within weeks in the atmosphere, each ensemble member evolves chaotically, affected by atmospheric circulation fluctuations characteristic of a random, stochastic process (e.g., Lorenz 1963; Deser et al. 2012b)”.. Each simulation is forced by repeating annual cycles of sea-surface temperature and insolation. The temperature, humidity, two wind components at each grid point, and selected vertical fluxes at the surface and top of the atmosphere in each grid column are saved every six hours.
For ML training, the temperature, humidity, and two wind components are averaged along FV3GFS’s 63 levels to 8 vertical layers, and the data are interpolated to a latitude-longitude grid of $180\times 360$ dimensions.
@@ Appendix C Implementation details
All methods and baselines are conditioned on the forcings, $\boldsymbol{f}_{t}$, by simple concatenation of the forcings with the remaining input variables across the channel dimension.
@@ C.1 Training and inference pseudocode
In Algorithm 1 we provide the procedure used to train our proposed method.
@@
In Algorithm 2 we provide the procedure used to sample with our proposed method.
@@
C.2 Hyperparameters
Architectural hyperparameters.
Table 4: Table is directly taken from [58], and reports the SFNO hyperparameters used for ACE as well as the interpolator and forecasting networks of our method.
@@ C.2 Hyperparameters
To fairly compare against the deterministic SFNO model from [58], we use exactly the same hyperparameters for training the interpolator and forecasting networks for our method, as described in Table 4555Names correspond to the definition of the SphericalFourierNeuralOperatorNet class found here: https://github.com/ai2cm/modulus/blob/94f62e1ce2083640829ec12d80b00619c40a47f8/modulus/models/sfno/sfnonet.py#L292. All configuration options not listed here are set to the defaults at the linked code.. For the stochastic version of ACE, ACE-STO, we re-train ACE from scratch with the only difference being that we use a dropout rate of $10\%$ for the MLP in the SFNO architecture. We train the stochastic interpolator model, $\text{SFNO}_{\phi}$, in our method using the same dropout rate. Both of these stochastic models are run using MC dropout (i.e. enabling the dropout layers at inference time).
Optimization hyperparameters.
Table 5: Optimization hyperparameters. The effective batch size is calculated as batch size $\times$ number of GPUs $\times$ accumulate_grad_batches, and is ensured to be the same for all our trained models regardless of the number of GPUs used. Gradient accumulation refers to the number of forward passes that are aggregated before a gradient step is taken by the optimizer.
@@
Learning rate schedule Cosine annealing (single cycle)
@@
We use the PyTorch Lightning framework to train our models. We train the interpolator networks for DYffusion and our method on the same relative L2 loss function used for the baseline from [58], and the corresponding forecaster networks on the L1 loss. The models that we train on our own, i.e. the interpolation and forecasting networks of DYffusion and our method are trained with mixed precision. Inference is always run at full precision. More optimization-related hyperparameters are discussed in Table 5.
Appendix D Additional results and figures
D.1 Additional ensemble metrics
Unless specified otherwise, all ensemble results are based on 25 ensemble members. We evaluate these ensembles with the ensemble-mean RMSE (eRMSE), the Continuous Ranked Probability Score (CRPS) [35], and the spread-skill ratio (spread / RMSE). The CRPS is a proper scoring rule and a popular metric in the probabilistic forecasting literature [16, 10, 44]. The spread-skill ratio is defined as the ratio of the square root of the ensemble variance to the corresponding ensemble-mean RMSE. It serves as a measure of the reliability of the ensemble, where values smaller than 1 indicate underdispersion (i.e. the probabilistic forecast is overconfident in its forecasts), and larger values overdispersion [14, 43]. All metrics are area-weighted according to the size of the grid cell. For the CRPS and RMSE lower is better, for spread / RMSE closer to 1 is better and indicates a well-dispersed ensemble.
@@ D.2 Climate Biases
We quantitatively analyze the 10-year time mean biases of our model and the baselines in terms of the global mean errors in Figure 6. The time-mean prediction is the average over the 14,600 predicted snapshots during the 10 years. Our method significantly reduces climate biases over the baseline methods across most fields. Notably, the errors of our method are often closer to the noise floor of the reference simulation than to the next best baseline. We also show that our method can be effectively ensembled to further reduce climate biases. The ensemble-mean prediction reliably improves time-mean scores across all fields. Interestingly, the stochastic version of ACE, ACE-STO, significantly underperforms the deterministic version. Similarly, the direct application DYffusion fails to match the deterministic ACE baseline, even after ensemble-averaging. This shows that MC dropout and DYffusion alone are not the reason for the encouraging performance of our method, but rather the holistic integration of all components, including MC dropout in our SFNO-based interpolator network.
D.2.1 Zonal time-means
@@
In this section, we analyze the absolute magnitudes of the simulated time-means by examining their zonal averages, where zonal refers to aggregation over the longitude dimension. We also report the standard deviation of the respective ensembles of time- and zonal-means for the reference and stochastic methods. We visualize these in Figure 7 and Figure 8. For several fields, including surface pressure, total water path (not shown), and near-surface temperature (top left subplot in Fig. 7), differences between the simulations are not visually noticeable, except for some biases near the poles by the baseline methods. Differences become more apparent as we analyze fields at higher altitude and wind fields. Generally, our method provides the closest emulation to the reference model. Although near-surface fields are the most relevant for society and decision-making, the clear biases of the baseline method at high-altitude levels might contribute to long-term biases, especially in longer simulations, due to the interactions of atmospheric dynamics across all levels. This observation may partly explain why our method achieves the lowest time-mean biases and RMSEs, as discussed in Section D.2.
@@
Figure 8: Zonal-means of the simulated 10-year time-mean climatologies for a representative subset of four northward (meridional) wind fields. Our method generally provides the closest emulation to the reference data, except the level-0 polar latitudes.
D.2.2 Climate variability

### v2

@@ B.3 FV3GFS configuration for data generation
In the following we summarize the reference data for this study, as also discussed in Section 2.1 of [67]. The training and validation data is generated by running an ensemble of 11 10-year (after discarding a 3-month spinup period) FV3GFS simulations on a C96 cubed-sphere grid (approximately 100 km horizontal grid spacing) with 63 vertical levels. The simulations are an initial-condition ensemble. That is, they are identical except for using different initial atmospheric states. Initial-condition ensembles are a popular tool in climate modeling [32]. Discarding a 3-month spinup period ensures that each simulation is independent of each other due to the chaoticity of the atmosphere [^7] E.g. see Kay et al. [32], who note that: “After initial condition memory is lost, which occurs within weeks in the atmosphere, each ensemble member evolves chaotically, affected by atmospheric circulation fluctuations characteristic of a random, stochastic process (e.g., Lorenz 1963; Deser et al. 2012b)”.. Each simulation is forced by repeating annual cycles of sea-surface temperature and insolation. The temperature, humidity, two wind components at each grid point, and selected vertical fluxes at the surface and top of the atmosphere in each grid column are saved every six hours. For ML training, the temperature, humidity, and two wind components are averaged along FV3GFS’s 63 levels to 8 vertical layers, and the data are interpolated to a latitude-longitude grid of $180\times 360$ dimensions.
@@ Appendix C Implementation details
All methods and baselines are conditioned on the forcings, $\boldsymbol{f}_{t}$, by simple concatenation of the forcings with the remaining input variables across the channel dimension. We use PyTorch Lightning [16] and Weights & Biases [6] as part of our software stack.
@@ C.1 Training and inference pseudocode
In Algorithms 1 and 2 we provide the procedures used to train and sample from our proposed method, respectively.
@@
C.2 Discussion on the training horizon
The training horizon, $h$, is a critical hyperparameter for both DYffusion and our proposed method. Throughout this study, we use $h=6$ (corresponding to 36 hours) for both approaches. While we initially explored other horizons, we chose $h=6$ as it strikes an optimal balance: A smaller horizon (e.g., $h=3$) reduces the number of sampling steps since the reverse sampling process directly corresponds to physical time steps, potentially degrading performance. Conversely, a larger horizon makes the forecasting task more challenging, as predicting $\boldsymbol{x}_{t+h}$ from $\boldsymbol{x}_{t}$ becomes increasingly difficult for the forecasting model.
Our choice is further supported by the DYffusion paper, which successfully used $h=7$ for sea surface temperature forecasting. While we believe that values close to $h=6$ would likely perform similarly well, comprehensive ablation studies would require re-training two neural networks sequentially, making such experiments computationally expensive to run.
C.3 Hyperparameters
Figure 7: Table is directly taken from [67], and reports the SFNO hyperparameters used for ACE as well as the interpolator and forecasting networks of our method.
@@ C.3 Hyperparameters
Figure 8: Optimization hyperparameters. The effective batch size is calculated as data loader batch size $\times$ number of GPUs $\times$ number of gradient accumulation steps, and is ensured to be the same for all our trained models regardless of the number of GPUs used.
@@
Learning rate schedule Cosine annealing
@@
Architectural hyperparameters.
To fairly compare against the deterministic SFNO model from [67], we use exactly the same hyperparameters for training the interpolator and forecasting networks for our method, as described in Table $8$Names correspond to the definition of the SphericalFourierNeuralOperatorNet class found at: https://github.com/ai2cm/modulus/blob/94f62e1ce2083640829ec12d80b00619c40a47f8/modulus/models/sfno/sfnonet.py#L292. Unless specified otherwise, defaults are used. . For the stochastic version of ACE, ACE-STO, we re-train ACE from scratch with the only difference being that we use a dropout rate of $10\%$ for the MLP in the SFNO architecture. We train the stochastic interpolator model, $\text{SFNO}_{\phi}$, in our method using the same dropout rate. Both of these stochastic models are run using MC dropout (i.e. enabling the dropout layers at inference time). For our interpolator network, we also use a $10\%$ rate for stochastic depth [28], which is also enabled at inference time. This choice was informed by preliminary experiments focused on training a good interpolator network. There, we found the addition of stochastic depth to slightly improve the interpolator’s validation CRPS scores (for the interpolated timesteps 1 to 5) and significantly improve the calibration of the interpolation ensemble based on the spread-skill ratio (averaged across variables from around 0.26 to 0.35). We found worse results when using stochastic depth for ACE-STO at inference time.
Optimization hyperparameters.
@@
We train the interpolator networks for DYffusion and our method on the same relative L2 loss function used for the baseline from [67], and the corresponding forecaster networks on the L1 loss. The models that we train on our own, i.e. the interpolation and forecasting networks of DYffusion and our method are trained with mixed precision. Inference is always run at full precision. For the non-interpolation networks, we perform early stopping based on the best CRPS averaged over a 500-step (125 days) rollout. More optimization-related hyperparameters are discussed in Table 8.
Appendix D Metrics
Unless specified otherwise, all ensemble results are based on $E=25$ ensemble members. All metrics are area-weighted according to the size of the grid cell, as described below.
D.1 Preliminaries
Let $\mathbf{X}\in\operatorname{\mathbb{R}}^{E\times I\times J}$ denote an ensemble of predictions, and $\mathbf{Y}\in\operatorname{\mathbb{R}}^{I\times J}$ the corresponding targets, where $E$ is the number of ensemble members, $I$ is the number of latitudes, and $J$ the number of longitudes in the grid. In the context of this paper, $\mathbf{Y}$ usually corresponds to the validation, reference 10-year time-mean and $\mathbf{X}$ corresponds to an ensemble of 10-year time-means simulated by the reference climate model (excluding the validation time-mean), our proposed method, or any of the baselines.
Let $w(i)$ denote the normalized latitude-dependent area weights at latitude $i$, such that $\frac{1}{I}\sum_{i}^{I}w(i)=1$, which ensure that spatial means are not biased towards the polar regions (see e.g. Rasp et al. [53]).
D.2 Member-wise Metrics
We report the average, member-wise area-weighted bias, Mean Absolute Error (MAE) and Root Mean Square Error (RMSE), which are defined as follows
Bias $=\frac{1}{EIJ}\sum_{e=1}^{E}\sum_{i,j}w(i)(\mathbf{X}_{e,i,j}-\mathbf{Y}_{i,j})$ (1)
MAE $=\frac{1}{EIJ}\sum_{e=1}^{E}w(i)\lvert\mathbf{X}_{e,i,j}-\mathbf{Y}_{i,j}\rvert$ (2)
RMSE $=\frac{1}{E}\sum_{e=1}^{E}\sqrt{\frac{1}{IJ}\sum_{i,j}w(i)(\mathbf{X}_{e,i,j}-\mathbf{Y}_{i,j})^{2}}$ (3)
For the bias, closer to zero is better, for the MAE and RMSE lower is better.
D.3 Ensemble Metrics
Ensemble-mean RMSE.
For a skillful ensemble, the magnitude of the average, member-wise RMSE (see above) can be reduced by computing the RMSE on the ensemble-mean prediction, defined as $\bar{\mathbf{X}}_{i,j}=\frac{1}{E}\sum_{e=1}^{E}\boldsymbol{x}_{e,i,j}$, instead.
$\text{RMSE}_{\text{ens}}=\sqrt{\frac{1}{IJ}\sum_{i,j}w(i)(\bar{\mathbf{X}}_{i,j}-\mathbf{Y}_{i,j})^{2}}$ (4)
Spread-Skill Ratio (SSR).
Following Fortin et al. [18], the spread-skill ratio is defined as the ratio between the ensemble spread and the ensemble-mean RMSE. The ensemble spread is defined as the the square root of the ensemble variance
$\text{Spread}=\sqrt{\frac{1}{IJ}\sum_{i,j}w(i)\text{var}_{e}(\mathbf{X}_{e,i,j})},$ (5)
where $\text{var}_{e}$ computes the variance of the ensemble.Then, we can compute the spread-skill ratio simply as
$\text{SSR}=\sqrt{\frac{E+1}{E}}\frac{\text{Spread}}{\text{RMSE}_{\text{ens}}},$ (6)
where $\sqrt{\frac{E+1}{E}}$ is a correction factor which is especially important to include for small ensemble sizes. Note that this factor is omitted in e.g. WeatherBench-2 [53]. The SSR serves as a simple measure of the reliability of the ensemble, where values smaller than 1 indicate underdispersion (i.e. the model is overconfident in its predictions), and larger values overdispersion. That is, closer to 1 is better.
Continuous Ranked Probability Score (CRPS).
Following Zamo and Naveau [72], we use the unbiased version of the CRPS [44], which is a proper scoring rule:
$\text{CRPS}=\frac{1}{IJ}\sum_{i,j}w(i)\left[\frac{1}{E}\sum_{e=1}^{E}|\mathbf{X}_{e,i,j}-\mathbf{Y}_{i,j}|-\frac{1}{2E(E-1)}\sum_{e=1}^{E}\sum_{f=1}^{E}|\mathbf{X}_{e,i,j}-\mathbf{X}_{f,i,j}|\right]$ (7)
where the first term represents the skill, and the second term represents the spread. The biased CRPS averages over the spread with the factor $\frac{1}{2E^{2}}$, which is biased–especially for small ensemble sizes–compared to using the unbiased version with the factor $\frac{1}{2E(E-1)}$. Note that common Python packages such as xskillscore and properscoring use the biased version. Lower is better.
Note on deterministic models.
For deterministic models like ACE without initial condition ensembling, the ensemble size is trivially $E=1$, causing $\bar{\mathbf{X}}_{i,j}$ to be identical to $\mathbf{X}$. This results in $\text{RMSE}_{\text{ens}}$ reducing to standard RMSE, MAE equaling CRPS, and a zero spread-skill ratio. To accurately reflect that ensemble metrics are not meaningful for single-member deterministic predictions, we denote these metrics as -- for ACE in Table 4. While incorporating initial condition ensembling would enhance ACE’s performance on these metrics beyond the naive deterministic baseline, such techniques are orthogonal to the model-based ensembling approaches explored in this work. We leave this extension to future research, noting that initial condition ensembling could potentially improve results for all models in our comparison, including the inherently stochastic ones.
Appendix E Additional results and figures
@@ E.1 Climate Biases
We quantitatively analyze the 10-year time mean biases of our model and the baselines in terms of the global mean RMSE in Figure 9. The time-mean prediction is the average over the 14,600 predicted snapshots during the 10 years. Our method significantly reduces climate biases over the baseline methods across most fields. Notably, the errors of our method are often closer to the noise floor of the reference simulation than to the next best baseline. We also show that our method can be effectively ensembled to further reduce climate biases, its ensemble-mean reliably improving time-mean scores across all fields. Interestingly, the stochastic version of ACE, ACE-STO, significantly underperforms the deterministic version. Similarly, the direct application DYffusion fails to match the deterministic ACE baseline, even after ensemble averaging. This shows that MC dropout and DYffusion alone are not the reason for the encouraging performance of our method, but rather the holistic integration of all components, including MC dropout in our SFNO-based interpolator network.
In Table 4, we report a comprehensive evaluation of the (ensemble) of 10-year time-means of each method for a subset of ten representative variables. We report the mean bias error, mean absolute error (MAE), root mean square error (RMSE), ensemble-mean RMSE, spread-skill ratio, and Continuous Ranked Probability Score (CRPS), which are rigorously defined in Appendix D.
Table 4: Comprehensive evaluation of simulated 10-year time-means. Bias, RMSE, and MAE represent average member-wise scores. For Bias (Spread-skill ratio; SSR) closer to 0 (1) is better. For the other metrics, lower is better, with relative changes from the reference shown in parentheses. See Appendix D for mathematical formulations and Table 3 for variable descriptions and units.
Variable Metric Reference Ours ACE ACE-STO DYffusion
$\mathrm{TWP}$ Bias 0.004 -0.043 0.021 0.017 0.686
RMSE 0.336 0.404 (+20%) 0.653 (+94%) 1.372 (+308%) 0.965 (+187%)
$\text{RMSE}_{\text{ens}}$ 0.249 0.327 (+31%) - 1.206 (+385%) 0.934 (+276%)
SSR 1.017 0.760 - 0.574 0.273
MAE 0.245 0.303 (+24%) 0.459 (+88%) 0.957 (+291%) 0.768 (+214%)
CRPS 0.125 0.178 (+43%) - 0.639 (+413%) 0.644 (+417%)
$p_{s}$ Bias 0.036 4.820 45.47 34.82 -126.4
RMSE 39.37 48.79 (+24%) 103.5 (+163%) 131.1 (+233%) 151.5 (+285%)
$\text{RMSE}_{\text{ens}}$ 31.50 39.59 (+26%) - 120.1 (+281%) 149.0 (+373%)
SSR 0.847 0.766 - 0.470 0.190
MAE 26.26 35.60 (+36%) 71.69 (+173%) 93.14 (+255%) 134.8 (+413%)
CRPS 14.44 21.91 (+52%) - 66.48 (+360%) 121.6 (+742%)
$T_{7}$ Bias 0.011 -0.049 0.121 0.369 0.311
RMSE 0.172 0.290 (+69%) 0.349 (+103%) 0.831 (+383%) 0.692 (+302%)
$\text{RMSE}_{\text{ens}}$ 0.124 0.267 (+114%) - 0.734 (+490%) 0.684 (+450%)
SSR 1.065 0.474 - 0.634 0.158
MAE 0.108 0.187 (+73%) 0.224 (+108%) 0.510 (+373%) 0.408 (+278%)
CRPS 0.054 0.132 (+147%) - 0.343 (+540%) 0.360 (+573%)
$T_{5}$ Bias 0.005 -0.068 0.079 0.173 0.377
RMSE 0.171 0.244 (+42%) 0.333 (+94%) 0.610 (+256%) 0.540 (+215%)
$\text{RMSE}_{\text{ens}}$ 0.132 0.211 (+60%) - 0.525 (+299%) 0.527 (+301%)
SSR 0.933 0.619 - 0.657 0.228
MAE 0.117 0.171 (+46%) 0.243 (+108%) 0.451 (+286%) 0.388 (+232%)
CRPS 0.060 0.110 (+84%) - 0.299 (+402%) 0.330 (+455%)
$T_{0}$ Bias 0.000 0.034 0.162 -0.127 0.517
RMSE 0.124 0.220 (+78%) 0.444 (+259%) 0.767 (+520%) 0.592 (+379%)
$\text{RMSE}_{\text{ens}}$ 0.074 0.202 (+174%) - 0.674 (+815%) 0.585 (+695%)
SSR 1.533 0.517 - 0.599 0.161
MAE 0.084 0.150 (+78%) 0.316 (+277%) 0.550 (+555%) 0.526 (+527%)
CRPS 0.034 0.102 (+200%) - 0.348 (+921%) 0.481 (+1312%)
$u_{7}$ Bias 0.012 0.038 -0.170 -0.023 0.077
RMSE 0.240 0.307 (+28%) 0.456 (+90%) 0.935 (+289%) 0.462 (+92%)
$\text{RMSE}_{\text{ens}}$ 0.178 0.239 (+35%) - 0.874 (+391%) 0.427 (+140%)
SSR 1.012 0.846 - 0.412 0.438
MAE 0.173 0.226 (+31%) 0.343 (+98%) 0.693 (+300%) 0.339 (+96%)
CRPS 0.087 0.129 (+48%) - 0.519 (+494%) 0.249 (+185%)
$v_{7}$ Bias 0.005 0.015 0.009 0.044 -0.067
RMSE 0.196 0.224 (+14.3%) 0.299 (+53%) 0.592 (+202%) 0.320 (+64%)
$\text{RMSE}_{\text{ens}}$ 0.152 0.178 (+17.0%) - 0.548 (+260%) 0.292 (+92%)
SSR 0.910 0.802 - 0.439 0.471
MAE 0.138 0.164 (+18.6%) 0.224 (+62%) 0.440 (+218%) 0.247 (+79%)
CRPS 0.072 0.094 (+30%) - 0.325 (+351%) 0.179 (+148%)
$\mathrm{WS}_{7}$ Bias 0.003 -0.053 -0.017 -0.080 -0.001
RMSE 0.243 0.303 (+24%) 0.437 (+79%) 0.886 (+264%) 0.450 (+85%)
$\text{RMSE}_{\text{ens}}$ 0.183 0.238 (+30%) - 0.823 (+349%) 0.415 (+126%)
SSR 0.976 0.830 - 0.430 0.445
MAE 0.175 0.224 (+28%) 0.331 (+89%) 0.659 (+277%) 0.334 (+91%)
CRPS 0.089 0.128 (+44%) - 0.488 (+449%) 0.244 (+175%)
$\mathrm{WS}_{5}$ Bias 0.022 0.058 -0.104 -0.036 -0.081
RMSE 0.324 0.398 (+23%) 0.626 (+93%) 1.128 (+248%) 0.591 (+82%)
$\text{RMSE}_{\text{ens}}$ 0.240 0.311 (+30%) - 1.030 (+329%) 0.543 (+126%)
SSR 1.013 0.837 - 0.475 0.452
MAE 0.248 0.311 (+25%) 0.492 (+98%) 0.878 (+254%) 0.456 (+84%)
CRPS 0.124 0.176 (+42%) - 0.636 (+412%) 0.329 (+165%)
$\mathrm{WS}_{0}$ Bias 0.151 -0.022 -0.167 0.854 1.642
RMSE 0.450 0.944 (+110%) 2.163 (+381%) 2.661 (+491%) 3.158 (+602%)
$\text{RMSE}_{\text{ens}}$ 0.307 0.887 (+189%) - 2.349 (+664%) 3.142 (+922%)
SSR 1.203 0.397 - 0.573 0.110
MAE 0.336 0.752 (+124%) 1.626 (+384%) 2.035 (+506%) 2.044 (+509%)
CRPS 0.152 0.589 (+287%) - 1.354 (+789%) 1.874 (+1131%)
E.1.1 Zonal time-means
@@
In this section, we analyze the absolute magnitudes of the simulated time-means by examining their zonal averages (aggregated over the longitude dimension). We also visualize the standard deviation of the respective ensembles of time- and zonal-means for the reference and stochastic methods. We visualize these in Figures 10 and 11. For several fields, including surface pressure, total water path (not shown), and near-surface temperature (top left subplot in Fig. 10), differences between the simulations are not visually noticeable, except for polar biases in baseline methods. However, discrepancies become pronounced in higher-altitude and wind fields, where our method generally achieves the closest agreement with the reference model. Although near-surface fields are the most relevant for society and decision-making, the clear biases of the baseline method at high-altitude levels might contribute to long-term biases, especially in longer simulations, due to the interactions of atmospheric dynamics across all levels. This observation may partly explain why our method achieves the lowest time-mean biases and RMSEs, as discussed in Section E.1.
@@
Figure 11: Zonal-means of the simulated 10-year time-mean climatologies for a representative subset of four northward (meridional) wind fields. Our method generally provides the closest emulation to the reference data, except for the level-0 polar latitudes.
E.1.2 Climate variability